OpenStudy (anonymous):

Suppose that a culture of bacteria has an initial population of n=100. If the population doubles every three days, determine the number of bacteria present after 30 days. How much time is required to reach 4250 in number?

OpenStudy (anonymous):

uhm can you explain using the malthus method i guess

sam (.sam.):

Is the answer correct?

sam (.sam.):

Do you have the answer?

OpenStudy (anonymous):

it says 16.23 days

sam (.sam.):

Geometric progression

sam (.sam.):

First term is, \[a _{1} = 100\] After 3 days, which is the fourth term, Use formula \[a _{n}=a _{1}r ^{n-1}\] \[a_{4}=100r ^{3} =200\] r=1.2599

sam (.sam.):

n= days, When n days, (use back the formula) \[a _{n}=100(1.2599)^{n-1}=4250\]

sam (.sam.):

n= 17.22

sam (.sam.):

Use natural logarithm to solve.

OpenStudy (anonymous):

okie dokie...i dont know if this helps but in differential equations i'm suppose to use like this half life equation y=e^kt+c